OpenStudy (anonymous):

Use the distributive property to simplify the expression below: −2(d − 4) + 3d Please explain how to do it totally lost with this! :)

OpenStudy (anonymous):

2 times d and then times 4 and then write like this -2d+8+3d and then you combine like terms ok 2d and 3d

OpenStudy (anonymous):

Thank you now I get it ! (:

OpenStudy (anonymous):

ok first you multiply -2 by d and -4 which gives you -2d+4 because two negatives multiplied together makes a positive. now you have -2d+4+3d. so combine your like terms which gives you d+4 or 1d+4. it doesn't really matter if you put the one or not

OpenStudy (anonymous):

umm -2d--8 + 3d so then two -- is a + therefore -2d + 8 + 3d = 8+ 3d - 2d hence = 8 + d